# Networked Communication for Mean-Field Games with Function Approximation and Empirical Mean-Field Estimation

### v2

@@ Abstract
Recent algorithms allow decentralised agents, possibly connected via a communication network, to learn equilibria in Mean-Field Games from a non-episodic run of the empirical system. However, these algorithms are for tabular settings: this computationally limits the size of agents’ observation space, meaning the algorithms cannot handle anything but small state spaces, nor generalise beyond policies depending only on the agent’s local state to so-called ‘population-dependent’ policies. We address this limitation by introducing function approximation to the existing setting, drawing on the Munchausen Online Mirror Descent method that has previously been employed only in finite-horizon, episodic, centralised settings. While this permits us to include the mean field in the observation for players’ policies, it is unrealistic to assume decentralised agents have access to this global information: we therefore also provide new algorithms allowing agents to locally estimate the global empirical distribution, and to improve this estimate via inter-agent communication. We show theoretically that exchanging policy information helps networked agents outperform both independent and even centralised agents in function-approximation settings. Our experiments demonstrate this happening empirically, by an even greater margin than in tabular settings, and show that the communication network allows decentralised agents to estimate the mean field for population-dependent policies.
1 Introduction
@@
The mean-field game (MFG) framework Lasry and Lions (2007); Huang et al. (2006) can be used to address the difficulty faced by multi-agent reinforcement learning (MARL) regarding computational scalability as the number of agents grows Yardim and He (2024); Zeng et al. (2024). It models a representative agent as interacting not with other individual agents in the population on a per-agent basis, but instead with a distribution over the other agents, called the mean field. The MFG framework analyses the limiting case when the population consists of an infinite number of symmetric and anonymous agents, that is, they have identical reward and transition functions which depend on the mean-field distribution rather than on the actions of specific other players. The solution to this game is the mean-field Nash equilibrium (MFNE), which can be used as an approximation for the Nash equilibrium (NE) in a finite-agent game (which is harder to solve in itself), with the error in the solution reducing as the number of agents N tends to infinity Saldi et al. (2018); Anahtarci et al. (2023); Yardim et al. (2024); Toumi et al. (2024); Hu and Zhang (2024); Chen et al. (2024). MFGs have thus been applied to a wide variety of real-world problems: see Laurière et al. (2022a) for examples.
@@
Recent works argue that classical algorithms for solving MFGs rely on assumptions and methods that are likely to be undesirable in real-world applications (e.g. swarm robotics or autonomous vehicles), emphasising that desirable qualities for practical MFG algorithms include: learning from the empirical distribution of $N$ agents (i.e. this distribution is generated only by the policies of the agents, rather than being updated by the algorithm itself or an external oracle/simulator); learning online from a single, non-episodic system run (i.e. similar to above, the population cannot be arbitrarily reset by an external controller); model-free learning; decentralisation; and fast practical convergence Yardim et al. (2023); Benjamin and Abate (2023). While these works address these desiderata, they do so only in settings in which the state and action spaces are small enough that the Q-function can be represented by a table, limiting their approaches’ scalability.
@@
Moreover, in those works, as in many others on MFGs, agents only observe their local state as input to their Q-function (which defines their policy). This is sufficient when the solved MFG is expected to have a stationary distribution (‘stationary MFGs’) Laurière et al. (2022a); Xie et al. (2021); Anahtarci et al. (2023); Zaman et al. (2023); Yardim et al. (2023); Benjamin and Abate (2023). However, in reality there are numerous reasons for which agents may benefit from being able to respond to the current distribution. Recent work has thus increasingly focused on these more general settings where it is necessary for agents to have so-called ‘master policies’ (a.k.a. population-dependent policies) which depend on both the mean-field distribution and their local state Cardaliaguet et al. (2015); Carmona et al. (2016); Perrin et al. (2022); Wu et al. (2024); Laurière et al. (2022a, b, a).
@@
The distribution is a large, high-dimensional observation object, taking a continuum of values. Therefore a population-dependent Q-function cannot be represented exactly in a table and must be approximated. To address these limitations while maintaining the desiderata for real-world applications given in recent works, we introduce function approximation to the MFG setting of decentralised agents learning online from a single, non-episodic run of the empirical system, allowing this setting to handle larger state spaces and to accept the mean-field distribution as an observation input.To overcome the difficulties of training non-linear approximators in this context, we use the so-called ‘Munchausen’ trick, introduced by Vieillard et al. (2020) for single-agent RL, and extended to MFGs by Laurière et al. (2022b), and to MFGs with population-dependent policies by Wu et al. (2024).
@@
We particularly explore this in the context of networked communication between decentralised agents Benjamin and Abate (2023). We demonstrate that communication brings two specific benefits over the purely independent setting (while also removing the undesirable assumption of a centralised learner, which in the real world may be unrealistic, a computational bottleneck and a vulnerable single point of failure). Firstly, when the Q-function is approximated rather than exact, some updates lead to better performing policies than others. Allowing networked agents to propagate better performing policies through the population (reminiscent of the use of fitness functions in distributed evolutionary algorithms Eiben and Smith (2015); Hart et al. (2015)) leads to faster learning than in the purely independent case and often even than in the centralised case, as we show both theoretically and empirically. Secondly, we argue that in the real world it is unrealistic to assume that decentralised agents, endowed with local state observations and limited (if any) communication radius, would be able to observe the global mean-field distribution and use it as input to their Q-functions / policy. We therefore further contribute two setting-dependent algorithms by which decentralised agents can estimate the global distribution from local observations, and further improve their estimates by communication with neighbours.
@@
We focus on ‘coordination’ games, where agents can increase their individual rewards by following the same strategy as others and therefore have an incentive to communicate policies, even if the MFG setting itself is technically non-cooperative. Thus our work can be applied to real-world problems in e.g. traffic signal control, formation control in swarm robotics, and consensus and synchronisation e.g. for sensor networks Soleimani et al. (2024). [^1] We further preempt concerns about the appropriateness of communication in competitive settings by wondering whether self-interested agents would be any more likely to want to obey a central learner as has usually been assumed. Moreover we show that self-interested communicating agents can obtain higher returns than independent agents even in non-coordination games (Fig. 5), indicating that they do have incentive to communicate. In summary, our contributions are:
@@
    • We use Munchausen RL for the first time in an infinite-horizon MFG context (for finite-horizon see Laurière et al. (2022b); Wu et al. (2024)).
    • This constitutes the first use of function approximation for solving MFGs from a single, non-episodic run of the empirical system (for tabular settings see Yardim et al. (2023); Benjamin and Abate (2023)).
@@
  • Rather than assuming that agents have access to this global knowledge as in prior works, we present two additional novel algorithms allowing decentralised agents to locally estimate the empirical distribution and to improve these estimates by inter-agent communication.
@@
  • We show theoretically that networked agents may learn faster than both centralised and independent agents in the function-approximation setting.
@@
  • We support this with extensive experiments, where our results showcase the two benefits of the decentralised communication scheme, which significantly outperforms both the independent and centralised settings.
@@
The paper is structured as follows. We give preliminaries in Sec. 2 and our core learning and policy improvement algorithm in Sec. 3. We present our mean-field estimation and communication algorithms in Sec. 4, theoretical results in Sec. 5 and experiments in Sec. 6. Sup. Mat. D contains an extended ‘Related work’ section.
2 Preliminaries
2.1 Mean-field games
@@
We use the following notation. $N$ is the number of agents in a population, with $\mathcal{S}$ and $\mathcal{A}$ representing the finite state and common action spaces, respectively. The set of probability measures on a finite set $\mathcal{X}$ is denoted $\Delta_{\mathcal{X}}$, and $\mathbf{e}_{x}\in\Delta_{\mathcal{X}}$ for $x\in\mathcal{X}$ is a one-hot vector with only the entry corresponding to $x$ set to 1, and all others set to 0. For time $t\geq 0$, $\hat{\mu}_{t}$ = $\frac{1}{N}\sum^{N}_{i=1}\sum_{s\in\mathcal{S}}$ $\mathbb{1}_{s^{i}_{t}=s}\mathbf{e}_{s}$in $\Delta_{\mathcal{S}}$ is a vector of length $|\mathcal{S}|$ denoting the empirical categorical state distribution of the $N$ agents at time $t$. For agent $i\in 1\dots N$, $i$’s policy at time $t$ depends on its observation $o^{i}_{t}$. We explore three different forms that this observation object can take:
@@
Definition 1 (N-player symmetric anonymous games).
@@
Definition 2 (Induced mean-field flow).
@@
Definition 3 (Mean-field discounted return).
@@
Definition 4 (Best-response (BR) policy).
A policy $\pi^{*}$ is a best response (BR) against the mean-field flow $\boldsymbol{\mu}$ if it maximises the discounted return $V(\cdot,\boldsymbol{\mu})$; the set of these policies is denoted $BR(\boldsymbol{\mu})$, i.e.
@@
Definition 5 (Mean-field Nash equilibrium (MFNE)).
@@
Previous works have shown that, in tabular settings, it is possible for a finite population of decentralised agents (each of which is permitted to have a distinct population-independent policy $\pi^{i}$) to learn the MFNE using only the empirical distribution $\hat{\mu}_{t}$, rather than the exactly calculated infinite flow $\boldsymbol{\mu}$ Yardim et al. (2023); Benjamin and Abate (2023). This MFNE may be the goal in itself, or it can in turn serve as an approximate NE for the harder-to-solve game involving the finite population. In this work we provide algorithms to perform this process in non-tabular and population-dependent settings, and demonstrate them empirically.
2.2 (Munchausen) Online Mirror Descent
Instead of performing the computationally expensive process of finding a $BR$ at each iteration, we can use a form of policy iteration for MFGs called Online Mirror Descent (OMD). This involves beginning with an initial policy $\pi_{0}$, and then at each iteration $k$, evaluating the current policy $\pi_{k}$ with respect to its induced mean-field flow $\boldsymbol{\mu}=I(\pi_{k})$ to compute its Q-function $Q_{k+1}$. To stabilise the learning process, we then use a weighted sum over this and past Q-functions, and set $\pi_{k+1}$ to be the softmax over this weighted sum, i.e. $\pi_{k+1}(\cdot|s,\mu)=softmax\left(\frac{1}{\tau_{q}}\sum^{k}_{\kappa=0}{Q_{\kappa}}(s,\mu,\cdot)\right)$. $\tau_{q}$ is a temperature parameter that scales the entropy in Munchausen RL Vieillard et al. (2020); note that this is a different temperature to the one agents use when selecting which communicated parameters to adopt, denoted $\tau^{comm}_{k}$ (Sec. 3.2).
If the Q-function is approximated non-linearly using neural networks, it is difficult to compute this weighted sum. The so-called ‘Munchausen trick’ addresses this by computing a single Q-function that mimics the weighted sum using implicit regularisation based on the Kullback-Leibler (KL) divergence between $\pi_{k}$ and $\pi_{k+1}$ Vieillard et al. (2020). Using this reparametrisation gives Munchausen OMD (MOMD), detailed further in Sec. 3.1 Laurière et al. (2022b); Wu et al. (2024). MOMD does not bias the MFNE, and has the same convergence guarantees as OMD Hadikhanloo (2017); Perolat et al. (2021); Wu et al. (2024).
2.3 Networks

### v3

@@ Abstract.
Recent algorithms allow decentralised agents, possibly connected via a communication network, to learn equilibria in mean-field games from a non-episodic run of the empirical system. However, these algorithms are for tabular settings: this computationally limits the size of agents’ observation space, meaning the algorithms cannot handle anything but small state spaces, nor generalise beyond policies depending only on the agent’s local state to so-called ‘population-dependent’ policies. We address this limitation by introducing function approximation to the existing setting, drawing on the Munchausen Online Mirror Descent method that has previously been employed only in finite-horizon, episodic, centralised settings. While this permits us to include the mean field in the observation for players’ policies, it is unrealistic to assume decentralised agents have access to this global information: we therefore also provide new algorithms allowing agents to locally estimate the global empirical distribution, and to improve this estimate via inter-agent communication. We prove theoretically that exchanging policy information helps networked agents outperform both independent and even centralised agents in function-approximation settings. Our experiments demonstrate this happening empirically, and show that the communication network allows decentralised agents to estimate the mean field for population-dependent policies.
Key words and phrases:
Mean-Field Games, Communication Networks, Reinforcement Learning
1. Introduction
@@
The mean-field game (MFG) framework (lasry2007mean; huangMFG) can be used to circumvent the difficulty faced by multi-agent reinforcement learning regarding computational scalability as the number of agents grows (yardim2024exploiting; zeng2024single). [^1] Preliminary versions of this paper have been accepted at: MARW at AAAI’25; GAAMAL at IJCAI’25; MALTA at AAAI’25; ALA at AAMAS’25; EMAS at AAMAS’25; GAIW at AAMAS’25; OptLearnMAS-25 at AAMAS’25; RLDM’25. It models a representative agent as interacting not with other individual agents in the population on a per-agent basis, but instead with a distribution over the other agents, called the mean field. The MFG framework analyses the limiting case when the population consists of an infinite number of symmetric and anonymous agents, that is, they have identical reward and transition functions which depend on the mean-field distribution rather than on the actions of specific other players. The solution to this game is the mean-field Nash equilibrium (MFNE), which can be used as an approximation for the Nash equilibrium (NE) in a finite-agent game (which is harder to solve in itself), with the error in the solution reducing as the number of agents N tends to infinity (finite_proofs; Anahtarci2020QLearningIR; yardim2024mean; TOUMI2024111420; hu2024mfoml; chen2024periodic). MFGs have thus been applied to a wide range of real-world problems: see survey_learningMFGs for examples.
@@
Recent works argue that classical algorithms for solving MFGs rely on assumptions and methods that are likely to be undesirable in real-world applications (e.g. swarm robotics, autonomous vehicles), emphasising that desirable qualities for practical MFG algorithms include: learning from the empirical distribution of $N$ agents (i.e. this distribution is generated only by the policies of the agents, rather than being updated by the algorithm itself or an external oracle/simulator); learning online from a single, non-episodic system run (i.e. similar to above, the population cannot be arbitrarily reset by an external controller); model-free learning; decentralised learning; and fast practical convergence (policy_mirror_independent; benjamin2024networked). While these works address these desiderata, they do so only in settings in which the state and action spaces are small enough that the Q-function can be represented by a table, limiting their approaches’ scalability.
@@
Moreover, in those works, as in many others on MFGs, agents only observe their local state as input to their Q-function (which defines their policy). This is sufficient when the solved MFG is expected to have a stationary distribution (‘stationary MFGs’) (survey_learningMFGs; one_that_may_sample; Anahtarci2020QLearningIR; zaman2023oraclefree; policy_mirror_independent; benjamin2024networked). However, in reality there are numerous reasons why agents may benefit from being able to respond to the current distribution (discussed further in Appx. B). Recent work has thus increasingly focused on these more general settings where it is necessary for agents to have so-called ‘master policies’ (a.k.a. population-dependent policies) which depend on both the mean-field distribution and their local state (cardaliaguet2015masterequationconvergenceproblem; carmona_common; perrin2022generalization; wu2024populationaware; survey_learningMFGs; scalable_deep).
@@
The distribution is a large, high-dimensional observation object, taking a continuum of values. Therefore a population-dependent Q-function cannot be represented in a table and must be approximated. To address these limitations while maintaining the desiderata for real-world applications given in recent works, we introduce function approximation to the MFG setting of decentralised agents learning online from a single, non-episodic run of the empirical system, allowing this setting to handle larger state spaces and to accept the mean-field distribution as an observation input. To overcome the difficulties of training non-linear approximators in this context, we use the so-called ‘Munchausen’ trick, introduced first for single-agent RL (NEURIPS2020_2c6a0bae), extended to MFGs (scalable_deep), and then to MFGs with population-dependent policies (wu2024populationaware).
We particularly explore this in the context of networked communication between decentralised agents. Almost all prior work relies on a centralised node to learn on behalf of all the agents. In this context ‘centralised’ does not imply global observability of the whole population’s actions - which would generally make computation infeasible given the complexity of the problem - but rather that learning is only conducted from the samples of a single representative agent, whose policy updates are assumed to be automatically pushed to the rest of the population by the central node. Therefore to reduce confusion, we sometimes refer to ‘central-agent learning’ instead of ‘centralised learning’ in contrast to prior works. More recent works have recognised that the assumption of a central learner might be be unrealistic in the real world, as well as a computational bottleneck and a vulnerable single point of failure (benjamin2024networked; policy_mirror_independent).
@@
We demonstrate that networked communication brings two specific benefits over the purely independent setting, while also removing the undesirable assumption of a central learner. Firstly, when the Q-function is approximated rather than exact, some updates lead to better performing policies than others. Allowing networked agents to propagate better performing policies through the population leads to faster learning than in the purely independent case and very often even than in the central-agent case, as we show both theoretically and empirically (this method is reminiscent of the use of fitness functions in distributed evolutionary algorithms (Eiben2015; survivability), and similarly of ‘population-based training’ (jaderberg2017population)). Secondly, we argue that in the real world it is unrealistic to assume that decentralised agents, endowed with local state observations and limited (if any) communication radius, would be able to observe the global mean-field distribution and use it as input to their Q-functions / policy. We therefore further contribute two setting-dependent algorithms by which decentralised agents can estimate the global distribution from local observations, and further improve their estimates by communication with neighbours.
@@
We focus on ‘coordination’ games, where agents can increase their individual rewards by following the same strategy as others and therefore have an incentive to communicate policies, even if the MFG setting itself is technically non-cooperative. Thus our work can be applied to real-world problems in e.g. traffic signal control, robotic swarm formation control, vehicle platooning, and consensus and synchronisation e.g. for sensor networks (10757965). [^2] We further preempt concerns about communication in competitive settings by wondering whether self-interested agents would be any more likely to want to obey a central learner as has usually been assumed. Moreover we show that self-interested communicating agents can obtain higher returns than independent agents even in non-coordination games (Fig. 6), indicating that they do have incentive to communicate. In summary, our contributions are:
@@
    • We use Munchausen RL for the first time in an infinite-horizon MFG context (cf. finite-horizon (scalable_deep; wu2024populationaware)).
    • This constitutes the first use of function approximation for solving MFGs from a single, non-episodic run of the empirical system (cf. tabular settings (policy_mirror_independent; benjamin2024networked)).
@@
  • Instead of assuming that agents have access to this global knowledge as in prior works, we present two additional novel algorithms allowing decentralised agents to locally estimate the empirical distribution and to improve these estimates by inter-agent communication.
@@
  • We prove theoretically that networked agents can learn faster than even central-agent populations in the function-approximation setting.
@@
  • We support this with extensive experiments, where our results demonstrate the two benefits of the decentralised communication scheme, which significantly outperforms both the independent and central-agent architectures.
@@
The paper is structured as follows. Related work is given in Sec. 2. We give preliminaries in Sec. 3 and our core learning and policy-improvement algorithm in Sec. 4. We present our mean-field estimation and communication algorithms in Sec. 5, theoretical results in Sec. 6 and experiments in Sec. 7.
2. Related work
We refer the reader to benjamin2024networked for detailed discussion around the setting of networked communication in MFGs, and to survey_learningMFGs for a broader survey of MFGs. Our work is most closely related to benjamin2024networked, which introduced networked communication to the infinite-horizon MFG setting. However, this work focuses only on tabular settings rather than using function approximation as in ours, and only addresses population-independent policies.
scalable_deep uses Munchausen Online Mirror Descent (MOMD), similar to our method for learning with neural networks, but in a different setting: their mean-field distribution is updated in an exact way and an oracle supplies a central learner with rewards and transitions for it to learn a population-independent policy, in a finite-horizon, episodic setting. wu2024populationaware uses MOMD to learn population-dependent policies, albeit also with a central-agent method that exactly updates the mean-field distribution in a finite-horizon episodic setting. perrin2022generalization learns population-dependent policies with function approximation in infinite-horizon settings like our own, but does so in a central-agent, two-timescale manner without using the empirical mean-field distribution. zhang2024stochastic also uses function approximation along a non-episodic path, but involves a generic central agent learning an estimate of the mean field rather than using an empirical population. Approaches that directly update an estimate of the mean field must be able to generate rewards from this arbitrary mean field, even if they otherwise claim to be oracle-free. They are thus inherently centralised algorithms and rely on strong assumptions that may not apply in real-world problems. Conversely, we are interested in practical convergence in online, deployed settings, where the reward is computed from the empirical finite population.
subjective_equilibria addresses decentralised learning from a continuous, non-episodic run of the empirical system using either full or compressed information about the mean field, but agents are assumed to receive this information directly, rather than estimating it locally as in the algorithm we now present. They also do not consider function approximation or inter-agent communication. In the closely related but distinct area of mean-field RL, pomfrl does estimate the empirical mean-field distribution from the local neighbourhood, however agents are seeking to estimate the mean action rather than the mean-field distribution over states as in our MFG setting. Their agents also do not have access to a communication network by which they can improve their estimates.
3. Preliminaries
3.1. Mean-field games
@@
We use the following notation. $N$ is the number of agents in a population, with $\mathcal{S}$ and $\mathcal{A}$ representing the finite state and common action spaces, respectively. The set of probability measures on a finite set $\mathcal{X}$ is denoted $\Delta_{\mathcal{X}}$, and $\mathbf{e}_{x}\in\Delta_{\mathcal{X}}$ for $x\in\mathcal{X}$ is a one-hot vector with only the entry corresponding to $x$ set to 1, and all others set to 0. For time $t\geq 0$, $\hat{\mu}_{t}$ = $\frac{1}{N}\sum^{N}_{i=1}\sum_{s\in\mathcal{S}}$ $\mathds{1}_{s^{i}_{t}=s}\mathbf{e}_{s}$ $\in$ $\Delta_{\mathcal{S}}$ is a vector of length $|\mathcal{S}|$ denoting the empirical categorical state distribution of the $N$ agents at time $t$. For agent $\in 1\dots N, ii$’s policy at time $t$ depends on its observation $o^{i}_{t}$. We explore three different forms that this observation object can take:
@@
Definition 3.1 (N-player symmetric anonymous games).
@@
Definition 3.2 (Induced mean-field flow).
@@
Definition 3.3 (Mean-field discounted return).
@@
Definition 3.4 (Best-response (BR) policy).
A policy $\pi^{*}$ is a best response (BR) against the mean-field flow $\boldsymbol{\mu}$ if it maximises the discounted return $V(\cdot,\boldsymbol{\mu})$; the set of these policies is denoted $BR(\boldsymbol{\mu})$:
@@
Definition 3.5 (Mean-field Nash equilibrium (MFNE)).
@@
Previous works have shown that, in tabular settings, it is possible for a finite population of decentralised agents (each of which is permitted to have a distinct population-independent policy $\pi^{i}$) to learn the MFNE using only the empirical distribution $\hat{\mu}_{t}$, rather than the exactly calculated infinite flow $\boldsymbol{\mu}$ (policy_mirror_independent; benjamin2024networked). This MFNE may be the goal in itself, or it can in turn serve as an approximate NE for the harder-to-solve game involving the finite population. In this work we provide algorithms to perform this process in non-tabular and population-dependent settings, and demonstrate them empirically.
3.2. (Munchausen) Online Mirror Descent
Instead of finding a $BR$ at each iteration, which is computationally expensive, we can use a form of policy iteration for MFGs called Online Mirror Descent (OMD). This begins with an initial policy $\pi_{0}$, and then at each iteration $k$, evaluating the current policy $\pi_{k}$ with respect to its induced mean-field flow $\boldsymbol{\mu}=I(\pi_{k})$ to compute its Q-function $Q_{k+1}$. To stabilise the learning process, we then use a weighted sum over this and past Q-functions, and set $\pi_{k+1}$ to be the softmax over this weighted sum, i.e. $\pi_{k+1}(\cdot|s,\mu)=softmax\left(\frac{1}{\tau_{q}}\sum^{k}_{\kappa=0}{Q_{\kappa}}(s,\mu,\cdot)\right)$. $\tau_{q}$ is a temperature parameter that scales the entropy in Munchausen RL (NEURIPS2020_2c6a0bae); note that this is a different temperature to the one agents use when selecting which communicated parameters to adopt, denoted $\tau^{comm}_{k}$ (Sec. 4.2).
If the Q-function is approximated non-linearly using neural networks, it is difficult to compute this weighted sum. The so-called ‘Munchausen trick’ addresses this by computing a single Q-function that mimics the weighted sum using implicit regularisation based on the Kullback-Leibler (KL) divergence between $\pi_{k}$ and $\pi_{k+1}$ (NEURIPS2020_2c6a0bae). Using this reparametrisation gives Munchausen OMD (MOMD), detailed further in Sec. 4.1 (scalable_deep; wu2024populationaware). MOMD does not bias the MFNE, and has the same convergence guarantees as OMD (hadikhanloo2017learninganonymousnonatomicgames; perolat2021scaling; wu2024populationaware).
3.3. Networks